how can 54/15 = 3 hours and 36 minutes?
Answer by Maths68(1474)

54/15=3hours + 36 minutes
Divide 36 by 60 to change minutes into hours
54/15=3hours + 36/60 hour
54/15=3 + 36/60
Taking LCM on right hand side
54/15=(180 + 36)/60
54/15=216/60
54/15=108/30
54/15=54/15